OpenStudy (anonymous):

needd help its a fraction 5/6y+1/4=-1/6y+1/2

OpenStudy (anonymous):

multiply by 6y and 4 so multiply by 24y and get : 20 + 6y = -4 + 12y 24 = 6y y = 4